OpenStudy (anonymous):

HELP!!! Write the direct variation of the equation if y varies directly with x and y = 45 when x = 9.

OpenStudy (anonymous):

Hmm do you really need help, this is exactly same as the previous one

OpenStudy (anonymous):

No it's not....

OpenStudy (saifoo.khan):

\[y = k * x\] Now find the value for k. \[45 = k*9\] \[k = \frac{45}{9}\]\[k=5\] now, insert that into original equation. \[y = k * x\]\[y = 5x\]

OpenStudy (anonymous):

Thanks! Can you help me with another?